What is the complement and supplement of pi over 10?
Found 2 solutions by flame8855, lwsshak3:
Answer by flame8855(424)

complement = pi/2 - pi/10=4pi/10= 2pi/5
supplement = pi-pi/10 = 9pi/10

Answer by lwsshak3(11628)

What is the complement and supplement of pi over 10?
**
let c=complement
let s= supplement
π/10+c=π/2
c=π/2-π/10
c=5π/10-π/10
c=4π/10
c=2π/5 or .4π
complement=.4π radians
..
π/10+s=π
c=π-π/10
c=10π/10-π/10
c=9π/10 or .9π
supplement=.9π radians